# SAM-guided Graph Cut for 3D Instance Segmentation

### v2

@@ 3.3 Deep graph cut
The edge weights calculated in Sec. 3.2 can be used to assess the affinity between two superpoints. While it’s possible to perform graph cut based directly on these weights, there is a drawback with this approach that can lead to suboptimal results: the calculation of each edge’s affinity is independent, lacking a global perspective, which maybe harmful for the final results. For instance, consider two objects composed of superpoints sets $A$ and $B$, interconnected by $n$ edges. Even if most of these edges are correctly predicted to have low affinity, the erroneous high affinity prediction of just one edge can result in the incorrect segmentation of objects $A$ and $B$ together. To improve the robustness of segmentation and reduce the occurrence of such situations, we feed the graph model into a graph neural network (GNN), which has a certain receptive field and can utilize the information of the surrounding nodes and edges.
@@
We utilize the GNN to predict affinity scores of graph edges, which are fed into the graph cut algorithm for 3D segmentation. To process the input graph, we design a GNN that consists of graph convolutional layers and fully connected layers. The graph model is first passed through graph convolutional layers to extract features for each vertex. Then, we concatenate features of two vertices with the corresponding edge score computed in Sec. 3.2, which are fed into fully connected layers to predict the affinity between two vertices.
When performing inference on a scene, we first utilize the trained GNN to predict the affinity scores of all edges in the graph, and then we segment the graph according to the affinity scores. Then we perform graph cut to cut the edges with low affinity score and then cluster the connected vertices as the final segmentation result, please refer to Sec. 4 for implementation details.
Pseudo labels generation.
@@
We propose a strategy for training the GNN without 3D ground truth annotation. To supervise the network, we first generate pseudo labels based on a 2D segmentation model. For pseudo-labels, the most ideal case would be to obtain the correct affinity for all edges, but this is unrealistic. In fact, while we require a high degree of accuracy for pseudo-labels, the completeness of these labels is a lesser priority. For this purpose, we use a 2D segmentation network, CropFormer [39], for this task. We first ran CropFormer on all views to obtain the instance segmentation results. For each pair of co-visible superpoints in every view, we record whether these superpoints are within the same mask. If they are co-visible in at least $n$ views and their records are consistent across all these n views, then we treat the pair as a pseudo-label. For example, if two superpoints are within the same mask in all co-visible views, they are treated as a positive sample in the pseudo-labels, vice versa.
@@
The reason for the choice is based on our empirical observation that CropFormer tends to yield relatively more complete and accurate masks for common object categories. For example, CropFormer consistently segments a chair entirely, whereas SAM sometimes segments parts of it, such as a single chair leg. Although CropFormer has its advantages, we opt for SAM in the previous graph construction stage due to the following considerations: Due to SAM’s unique prompt mechanism, it can use superpoints to control the granularity of segmentation to some extent. Moreover, SAM’s design, which allows for predicting multiple masks from a single prompt, makes it more adept at handling uncommon objects. Consequently, we chose SAM for constructing the entire graph, while using CropFormer to generate relatively incomplete but high-quality pseudo-labels.
Figure 4: 3D segmentation results on ScanNet, ScanNet++ and KITTI-360 datasets. Please zoom in for details. Compared to Mask3D, our method exhibits significantly better generalization on ScanNet++ and KITTI-360 datasets. Moreover, in comparison to SAM3D, our approach can segment objects in the scene more completely and accurately. We observed that Panoptic Lifting struggles to extract satisfactory geometry, so we leave the qualitative comparison with it to the supplementary material.
Training of the GNN.
@@
We employ pseudo-labels generated by CropFormer as supervision. For the edges included in these pseudo-labels, we compare the affinity scores of these edges predicted by GNN with the labels to binary cross-entropy loss, which is defined as
@@
where $s$ is the predicted affinity score and $s_{\text{ps}}$ is the pseudo-label of corresponding edge. Since pseudo-labels are sparse, we observed that direct training in this manner tends to limit the network’s capabilities. To improve the accuracy of the graph network’s predictions, we introduce an additional regularization for edges not included in the pseudo-labels during training. This regularization aims to ensure the predicted affinity score $s$, and the edge weight predicted by SAM, $w_{\text{SAM}}$, to be as consistent as possible. This loss is defined as
@@
With this design, the closer $w_{\text{SAM}}$ is to 0 or 1, the greater the penalty for inconsistency between $s$ and $w_{\text{SAM}}$.
@@ 4 Implementation details
When sampling points within the projection of a superpoint as prompts for SAM, we uniformly sample $k=5$ points within the projected mask. Considering that there might be slight inaccuracies in the camera poses, the points too close to the boundaries of projection masks could potentially fall outside the object. Therefore, we take care to avoid sampling these points during this process.
ScanNet ScanNet++ KITTI-360
@@
Felzenswalb 1.9 5.8 22.9 5.8 12.4 29.7 - - -
Guinard 0.6 2.4 24.0 0.8 2.8 16.5 2.1 6.2 24.1
SAM3D (w/o ensemble) 6.3 17.9 47.3 1.7 5.4 20.9 2.5 7.5 22.4
SAM3D (w/ ensemble) 13.7 29.7 54.5 8.3 17.5 33.7 6.3 16.0 35.6
SAM-based Graph + NCuts 5.1 13.6 39.5 7.5 16.5 34.2 10.1 18.4 31.0
SAM-based Graph + DBSCAN 6.6 14.1 24.7 8.2 15.0 22.6 12.9 23.5 35.6
SAM-based Graph + Graph partition 12.9 30.0 57.7 12.6 24.5 40.2 12.6 25.5 41.9
Ours 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2
Table 1: Quantitative results of 3D segmentation on ScanNet, ScanNet++ and KITTI-360 datasets. We report the AP scores averaged on all test scenes. Our method significantly outperforms the baseline methods on all datasets.
Training and inference of GNN.
We implement the GNN with PyTorch [35] and PyG [15]. When generating pseudo labels with CropFormer, we only consider the superpoint pairs that are co-visible in at least $n=10$ views with all consistent records. We train the GNN on ScanNet for 200 epochs, which takes about 20 minutes on NVIDIA A6000 GPU.
When performing segmentation based on the output of GNN, for every two vertices, if their affinity is below a certain threshold, we consider them to be unconnected. Conversely, if their affinity is above this threshold, to further improve robustness, we identify paths of length 2 between these two vertices. We then record the number of paths where both edges have high affinity scores and the number where one edge is high and the other low. If the ratio of the latter exceeds a predefined threshold, we regard the two vertices to be unconnected; otherwise, they are considered as connected. Once the connection status of each edge is determined using this method, we employ a union-find algorithm [46] to merge all connected superpoints, resulting in the 3D segmentation of the scene.
5 Experiments
5.1 Datasets, metrics and baselines
Datasets.
We perform the experiments on ScanNet (V2) [6], ScanNet++ [59] and KITTI-360 [30]. ScanNet is a large-scale RGB-D dataset that contains 1613 indoor scenes with geometry acquired by BundleFusion [7] and images captured by iPad Air2. ScanNet++ contains 280 indoor scenes with high-fidelity geometry acquired by the Faro Focus Premium laser scanner as well as high-resolution RGB images captured by iPhone 13 Pro and a DSLR camera with a fisheye lens. KITTI-360 is a large outdoor dataset with 300 suburban scenes, which consists of 320k images and 100k laser scans obtained through a mobile platform in a driving distance of 73.7km. Both these datasets are annotated with ground truth camera poses and instance-level semantic segmentations. We train the GNN on 1201 training scenes of ScanNet, we then evaluate our method on 312 validation scenes of ScanNet, 50 of ScanNet++ and 61 scenes of KITTI-360, according to the official split of each dataset.
Metrics.
[…]
ScanNet Mask3D 26.9 44.4 57.5
Ours 19.0 38.9 59.9
ScanNet++ Mask3D 8.8 15.0 22.3
Ours 14.3 26.7 43.5
KITTI-360 Mask3D 0.1 0.4 4.2
Ours 14.7 28.0 43.2
Table 2: Comparison with Mask3D. While Mask3D shows slightly better performance than ours on ScanNet, when comparing generalization results on ScanNet++ and KITTI-360, our method significantly outperforms Mask3D.
Baselines.
We compare our method with the following baselines: (1) Traditional segmentation method: [14, 17], which use purely geometric information to perform segmentation. (2) 2D-to-3D lifting method: SAM3D [58] and Panoptic Lifting [44]. We report the results of SAM3D with and without ensemble process. Since Panoptic Lifting is based on NeRF [33] and requires hours for per-scene optimization, we only report corresponding qualitative analyses. (3) Point cloud segmentation method: Mask3D [42], we use their official pretrained model which is trained on ScanNet dataset. (4) We segment our SAM annotated graph (without GNN) with traditional spectral clustering methods: Normalized Cuts [43] and DBSCAN [41], as well as the graph cut used in our method.
5.2 Comparisons with the state-of-the-art methods
We evaluate 3D segmentation metrics on ScanNet, ScanNet++ and KITTI-360 datasets. Averaged quantitative results are shown in Tab. 1. We also provide qualitative results in Fig. 4. By analysing quantitative and qualitative results, we found that our method significantly outperforms state-of-the-art unsupervised methods.
Panoptic Lifting can achieve reasonably good segmentation results in simpler scenes, but its performance deteriorates in more complex environments with a large number of objects. SAM3D can handle complex scenes, however, its segmentation often results in both large structures, like floors, and smaller objects, like chairs, being divided into multiple segments. In contrast, our method is able to produce more accurate and complete segmentation results. In our experiments with Normalized Cuts and DBSCAN, we found that carefully tuning hyperparameters can yield good result for a single scene. However, each scene varies significantly in scale and the number of objects, leading to considerable differences in the optimal hyperparameters. When we apply uniform hyperparameters across all dataset, the averaged metrics are not ideal.
Comparisons with Mask3D
In addition to unsupervised baselines, we also compare our method with Mask3D, which is the state-of-the-art supervised learning method. When evaluation, we use the official pretrained model provided by Mask3D, which is trained on ScanNet dataset with GT annotations. For fair comparison, we ignore floor and wall regions when evaluating on ScanNet and ScanNet++ datasets. The quantitative results are shown in Tab. 2. The results show that our method achieve much better generalization performance.
[…]
w/o GNN 12.9 30.0 57.7
w/o node features 13.4 30.7 58.0
w/o edge weights 1.9 5.6 22.0
w/o regularization loss 9.1 20.2 39.9
our full method 15.1 33.3 59.1
Table 3: Ablation studies on ScanNet dataset.

### v3

@@ 3.3 Deep graph cut
We first propose a simple segmentation method that is not based on neural networks. Specifically, based on the superpoint graph constructed in Sec. 3.1 and the edge weights of the graph calculated in Sec. 3.2, we use the edge weights to determine whether two superpoints are connected, then employ a method based on the union-find algorithm [52] to merge all connected superpoints, thereby achieving 3D segmentation of the scene.
@@
To improve the robustness of segmentation, we feed the graph model into a graph neural network (GNN) before segmentation. The GNN has a certain receptive field and can utilize the information of the surrounding nodes and edges and can predict edge affinity scores which can be more reliable than original edge weights. To process the input graph, we design a GNN that consists of graph convolutional layers and fully connected layers. The graph model is first passed through graph convolutional layers to extract features for each vertex. Then, we concatenate features of two vertices with the corresponding edge weight computed in Sec. 3.2, which are fed into fully connected layers to predict the affinity between two vertices. After that, we apply the same segmentation method based on the predicted affinity scores.
@@
Pseudo labels generation.  We propose a strategy for training the GNN without 3D ground truth annotation. To supervise the network, we first generate pseudo labels based on a 2D segmentation model. For pseudo-labels, the most ideal case would be to obtain the correct affinity for all edges, but this is unrealistic. In fact, while we require a high degree of accuracy for pseudo-labels, the completeness of these labels is a lesser priority. For this purpose, we use a 2D segmentation network, CropFormer [41], for this task. We first ran CropFormer on all views to obtain the instance segmentation results. For each pair of co-visible superpoints in every view, we record whether these superpoints are within the same mask. If they are co-visible in at least $n$ views and their records are consistent across all these $n$ views, then we treat the pair as a pseudo-label. For example, if two superpoints are within the same mask in all co-visible views, they are treated as a positive sample in the pseudo-labels, vice versa.
@@
The reason for choosing CropFormer is based on our empirical observation that it tends to yield relatively more complete and accurate masks for common object categories. For example, CropFormer consistently segments a chair entirely, whereas SAM sometimes segments parts of it, such as a single chair leg. Although CropFormer has its advantages, we opt for SAM in the previous graph construction stage due to the following considerations: SAM’s unique prompt mechanism can use superpoints to control the granularity of segmentation to some extent. Moreover, SAM’s design, which allows for predicting multiple masks from a single prompt, makes it more adept at handling uncommon objects. Consequently, we chose SAM for constructing the entire graph, while using CropFormer to generate relatively incomplete but high-quality pseudo-labels.
@@
Training of the GNN.  We employ pseudo-labels generated by CropFormer as supervision. For the edges included in these pseudo-labels, we compare the affinity scores of these edges predicted by GNN with the labels to calculate binary cross-entropy loss, which is defined as
@@
where $s$ is the predicted affinity score and $s_{\text{ps}}$ is the pseudo-label of corresponding edge. Since pseudo-labels are sparse, we observed that direct training in this manner tends to limit the network’s capabilities. To improve the accuracy of the graph network’s predictions, we introduce an additional regularization for edges not included in the pseudo-labels during training. This regularization aims to ensure the predicted affinity score $s$, and the corresponding edge weight predicted by SAM, $w_{\text{SAM}}$, to be as consistent as possible. This loss is defined as
@@
With this design, the closer $w_{\text{SAM}}$ is to 0 or 1, the greater the penalty for inconsistency between $s$ and $w_{\text{SAM}}$. The final loss is defined as $L=L_{\text{BCE}}+L_{\text{reg}}$.
@@ 4 Implementation details
Graph construction.  When sampling points within the projection of a superpoint as prompts for SAM, we uniformly sample $k=5$ points within the projected mask. Considering that there might be slight inaccuracies in the camera poses, the points too close to the boundaries of projection masks could potentially fall outside the object. Therefore, we take care to avoid sampling these points during this process.
Training of GNN.  We implement the GNN with PyTorch [37] and PyG [17]. When generating pseudo labels with CropFormer, we only consider the superpoint pairs that are co-visible in at least $n=10$ views with all consistent records. We train the GNN on ScanNet200 with pseudo-labels for 200 epochs, which takes about 20 minutes on an NVIDIA A6000 GPU.
5 Experiments
Figure 4: 3D segmentation results on ScanNet200, ScanNet++ and KITTI-360 datasets. Please zoom in for details. Compared to Mask3D, our method exhibits significantly better generalization on ScanNet++ and KITTI-360 datasets. Moreover, in comparison to SAM3D, our approach can segment objects in the scene more completely and accurately. We observed that Panoptic Lifting struggles to extract satisfactory geometry, so we leave the qualitative comparison with it to the supplementary material.
Table 1: Quantitative results of 3D segmentation on ScanNet200, ScanNet++ and KITTI-360 datasets. We report the AP scores averaged on all test scenes. Our method significantly outperforms the baseline methods on all datasets.
ScanNet200 ScanNet++ KITTI-360
mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25
Felzenswalb 4.8 9.8 27.5 8.8 16.9 36.1 - - -
Guinard 2.9 8.2 33.1 4.3 10.6 32.3 9.3 18.9 39.6
SAM3D (w/o ensemble) 12.1 28.6 54.1 3.0 7.9 22.3 4.6 10.6 26.0
SAM3D (w/ ensemble) 20.9 34.8 51.4 9.3 16.6 29.5 13.0 24.2 41.1
Ours + NCuts 15.7 31.7 59.0 10.1 18.6 34.7 19.5 30.2 45.0
Ours + DBSCAN 10.3 18.6 27.8 10.5 17.2 25.0 20.5 31.4 42.1
Ours (w/o GNN) 19.7 37.7 61.6 13.7 25.2 43.0 22.6 36.2 48.5
Ours (w/ GNN) 22.1 41.7 62.8 15.3 27.2 44.3 23.8 37.2 49.1
5.1 Datasets, metrics and baselines
Datasets.  We perform the experiments on ScanNet200 [45], ScanNet++ [66] and KITTI-360 [32]. ScanNet200 is built on ScanNet [8], which is a large-scale RGB-D dataset that contains 1613 indoor scenes with geometry acquired by BundleFusion [9] and images captured by iPad Air2. ScanNet++ contains 280 indoor scenes with high-fidelity geometry acquired by the Faro Focus Premium laser scanner as well as high-resolution RGB images captured by iPhone 13 Pro and a DSLR camera with a fisheye lens. KITTI-360 is a large outdoor dataset with 300 suburban scenes, which consists of 320k images and 100k laser scans obtained through a mobile platform in a driving distance of 73.7 km. All of these datasets are annotated with ground truth camera poses and instance-level semantic segmentations. The GNN in our method is trained on 1201 training scenes of ScanNet200 with our generated pseudo-labels, we then evaluate our method on 312 validation scenes of ScanNet200, 50 of ScanNet++ and 61 scenes of KITTI-360, according to the official split of each dataset.
Metrics.  We evaluate our segmentation performance with the widely-used Average Precision (AP) score. We follow the standard defined in [42, 8, 45] for evaluation, calculating AP with thresholds of $50\%$ and $25\%$ (denoted as AP50 and AP25, respectively) as well as AP averaged with IoU thresholds from $50\%$ to $95\%$ with a step size of $5\%$ (mAP). Since our method as well as most of baseline methods are class-agnostic, we do not consider semantic class label in evaluation, which follows the setting of [46]. Additionally, we exclude the predicted instances in unannotated regions for all methods to facilitate a fairer comparison.
Baselines.  We compare our method with the following baselines: (1) Traditional segmentation methods: [16, 19], which only use geometric information to perform segmentation. (2) 2D-to-3D lifting method: SAM3D [65] and Panoptic Lifting [50]. We report the results of SAM3D with and without the ensemble process. Since Panoptic Lifting is based on NeRF [35] and requires hours for per-scene optimization, we only report corresponding qualitative analyses. (3) Point cloud segmentation method: Mask3D [48], we use their official pretrained models. (4) We segment our SAM annotated graph (without GNN) with traditional spectral clustering methods: Normalized Cuts [49] and DBSCAN [47], as well as the graph cut used in our method.
5.2 Comparisons with the state-of-the-art methods
We evaluate 3D segmentation metrics on ScanNet200, ScanNet++ and KITTI-360 datasets. Averaged quantitative results are shown in Tab. 1. We also provide qualitative results in Fig. 4. By analyzing quantitative and qualitative results, we found that our method significantly outperforms state-of-the-art unsupervised methods.
Panoptic Lifting can achieve reasonably good segmentation results in simpler scenes, but its performance deteriorates in more complex environments with a large number of objects. SAM3D can handle complex scenes, however, its segmentation often results in both large structures, like floors, and smaller objects, like chairs, being divided into multiple segments. In contrast, our method is able to produce more accurate and complete segmentation results. In the experiments of our method combined with Normalized Cuts and DBSCAN, we found that carefully tuning hyperparameters can yield relatively good result for a single scene. However, each scene varies significantly in scale and the number of objects, leading to considerable differences in the optimal hyperparameters. When we apply uniform hyperparameters across all dataset, the averaged metrics are not ideal.
Table 2: Comparison with Mask3D. While Mask3D shows better performance than ours on ScanNet200, it cannot generalize well to ScanNet++ and KITTI-360.
@@ 5.2 Comparisons with the state-of-the-art methods
ScanNet200 Mask3D GT of ScanNet200 53.3 71.9 81.6
Mask3D GT of ScanNetV2 45.1 62.6 70.5
Ours (w/o GNN) - 19.7 37.7 61.6
Ours (w/ GNN) Pseudo-GT of ScanNet200 22.1 41.7 62.8
ScanNet++ Mask3D GT of ScanNet200 4.6 10.5 22.9
Mask3D GT of ScanNetV2 3.7 7.9 15.6
Ours (w/o GNN) - 13.7 25.2 43.0
Ours (w/ GNN) Pseudo-GT of ScanNet200 15.3 27.2 44.3
KITTI-360 Mask3D GT of ScanNet200 0.2 0.9 7.0
Mask3D GT of ScanNetV2 0.3 1.0 8.0
Ours (w/o GNN) - 22.6 36.2 48.5
Ours (w/ GNN) Pseudo-GT of ScanNet200 23.8 37.2 49.1
Comparisons with supervised method.  In addition to unsupervised baselines, we also compare our method with Mask3D, which is the state-of-the-art supervised learning method. We evaluate the class-agnostic AP scores, and the results are shown in Tab. 2. When evaluation, we use their official pretrained models. Specifically, we first evaluate their model trained on the ScanNet200 training set, which shows very good results on the ScanNet200 validation set. However, when applied to the ScanNet++ dataset, there is a significant performance drop. This is because ScanNet++, although also indoor scene data, has a different data collection method from ScanNet200, indicating that the Mask3D method is quite sensitive to differences in aspects such as the point cloud collection method. When applied to KITTI-360, the performance becomes extremely poor, as KITTI-360 is an outdoor scene, which is quite different from ScanNet200 so that Mask3D cannot generalize well. Then we evaluate Mask3D model trained on the ScanNetV2 training set. Compared to ScanNet200, ScanNetV2 lacks annotations for some categories of objects, resulting in decreased performance on the ScanNet200 validation set. This indicates that Mask3D has limited ability to generalize to different categories of objects even in the same type of scenes, and it also suffers significant performance declines on ScanNet++ and KITTI-360.
In contrast, our method (without GNN) achieves good results on all three datasets without any training, although our method performs worse on ScanNet200 than Mask3D, which is because Mask3D is trained with GT annotations provided by ScanNet200, allowing it to learn strong priors of this dataset. Our method shows a clear advantage on the ScanNet++ and KITTI-360 datasets. After using the GNN trained with pseudo-labels on ScanNet200, our method not only improves on the ScanNet200 validation set but also shows improvement on ScanNet++ and KITTI-360. This indicates that the GNN module in our method helps with segmentation and learns general prior for segmentation agnostic to the training dataset, thus has good generalization capabilities to data with large differences.
5.3 Ablation studies
To analyze the effectiveness of each module and design in our method, as well as the impact of hyper-parameters on performance, we conduct ablation studies on ScanNet200.
Table 3: Ablation studies of graph cut implementation.
mAP AP50 AP25
w/o GNN 19.7 37.7 61.6
w/o node features 19.4 37.5 61.3
w/o edge weights 10.1 21.5 42.2
w/o regularization loss 19.9 38.1 61.4
our full method 22.1 41.7 62.8
Table 4: Ablation studies of $k$.
$k$ mAP AP50 AP25
1 20.3 37.7 60.7
3 21.2 38.8 61.2
5 22.1 41.7 62.8
7 22.0 40.1 62.1
9 21.9 39.9 62.3
Table 5: Ablation studies of $n$.
$n$ mAP AP50 AP25
1 20.3 38.7 61.2
5 21.7 40.9 63.6
10 22.1 41.7 62.8
20 20.4 39.1 62.6
30 20.2 39.3 61.9
Ablation studies of graph cut implementation.  We evaluate with five configurations: (1) Apply graph cut directly on edge weights of the SAM annotated graph (without GNN refinement). (2) Annotate the graph without node features. (3) Annotate the graph without edge weights. (4) Train GNN without regularization loss. (5) Our full method. We report the quantitative results in Tab. 3. The results indicate that the absence of a GNN leads to a decline in segmentation performance. Node features have a certain impact on the predictive performance of the GNN, while edge weights have a very significant impact, almost playing a dominant role. Additionally, we found that regularization loss is also important.
Ablation studies of number of prompt points.  We conduct ablation studies on the number of prompt points $k$ sampled in each superpoint projection, the results are shown in Sec. 5.3. In our experiment, we chose $k=5$, and the ablation results suggest that our method is not very sensitive to the value of $k$, but it is affected to some extent. The results are relatively poor when $k=1$, as sampling only one point within a projection mask leads to significant information loss, especially when the projection mask area is large. When $k$ is larger, the performance slightly deteriorates, which may be due to SAM not being suitable for accepting too many points as a prompt.
Ablation studies of minimum number of co-visible views.  We conduct ablation studies on the minimum number of co-visible views $n$ required for generating pseudo labels, the results are shown in Sec. 5.3. In our experiment, we chose $n=10$, and the ablation results indicate that our method is robust to the choice of $n$, but our performance is relatively poorer when $n$ is either too small or too large. This is because when $n$ is small, a large number of unreliable pseudo-labels are generated. When $n$ is large, although the quality of the pseudo-labels is high, they become very sparse. In both cases, this leads to suboptimal results.
Ablation studies of pseudo-label generation.  Furthermore, we conduct experiments to analyze the advantages of using CropFormer to generate pseudo-labels compared to SAM. For the labels generated by CropFormer, we also use SAM to predict a label. The accuracy scores of both methods are evaluated using the ground truth annotations provided by ScanNet200. As indicated in the Tab. 6, CropFormer achieves higher accuracy.
6 Conclusion
In this paper, we introduced a novel 3D segmentation method with SAM guided graph cut. The key idea is to pre-segment 3D scenes into superpoints, and then utilize the prompt mechanism of SAM to assess the affinity scores between superpoints. We propose a GNN based graph cut method to achieve robust segmentation, which is trained with pseudo-labels generated by a 2D segmentation network. Experiments showed that the proposed method is able to achieve accurate segmentation results and can generalize well to different datasets.
Table 6: Analyses of the choice for pseudo-labels generation.